The sum of Andre and Butch's scores on a math test was 180. Andres score was two times as much as 30 less than Butch's score. Write an equation and find each student score
Answer by addingup(3677)

A+B = 180
A = 2(B-30)
A = 2B-30
--------------
A+B = 180
2B-30+B = 180
3B = 210
B = 70 Butch's score was 70 and
A = 2(70-30) = 80 Andre's score was 80